OpenStudy (anonymous):

need help solving these system of equations x-4y=33 4x+7y=-6

OpenStudy (anonymous):

so we need to solve for a variable, and to do that we should use substitution

OpenStudy (anonymous):

start with the top equation

OpenStudy (anonymous):

x-4y=33 -4y=33-x y=-33-x/4

OpenStudy (anonymous):

then substitute that y value into the second equation, to solve for x

OpenStudy (anonymous):

return to the first equation and use that x to solve for y

OpenStudy (anonymous):

I noticed you only put -4 is it written out like this x-x-4y=33

OpenStudy (anonymous):

and subtract x from 33

OpenStudy (anonymous):

-(33-x/4)

OpenStudy (anonymous):

\[-\frac{ 33-x }{ 4 }\]

OpenStudy (anonymous):

I kept writing it but thought I had it wrong

OpenStudy (anonymous):

mind your negatives

OpenStudy (anonymous):

thats the most frequent mistake

OpenStudy (anonymous):

yea that's what I always have a problem with

OpenStudy (anonymous):

can you feel your algebra muscles working ^^)

OpenStudy (anonymous):